A florist is paid $12 per hour and her helper is paid $8 per hour. Together they worked 30 hours on a job and made a total of $320. How many hours did the florist work?
Answer by josgarithmetic(39617)

This is like a mixture problem. Let x amount hours of the helper and y the hours of the florist.

8*x+12*y=320 accounting for dollars.
x+y=30 account for hours.

Divide the dollars equation by 4.
2*x+3*y=80

The system is as simple as
2*x+3*y=80
x+y=30
Subtract 2 times the second equation from the first equation:
0*x+1*y=20
y=20 hours for the florist.